The bond enthalpies of gas H2 , Cl2 and HCL are 104, 58 ,103 bar3 k-1 mol-1 .calculate the enthalpy of formation of HCl

Grade:11

1 Answers

Arun
6 years ago
1. The chemical equation for the formation of HCl is :

      H2 + Cl2 =2 HCl
2. Bond dissociation energy = Bonds broken - Bonds formed
                                            = (104 + 58) - 2 (103) = - 44 Kcal.